the length of a rectangle is 5 inches more than 3 times the width. If the perimeter is 26 inches, find the length and width

Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Length = 5 + 3x
Perimeter = 2x Length + 2x Width
26 = 2(5 + 3x) + 2(x)
26 = 10 + 6x + 2x
Collect like terms
2x + 6x = 26 - 10
8x = 16
x = 2inches.
Width = 2 inches
Length = 11 inches.
Hope this helps:-)